# Sound waves from primordial black hole formations

### v1

@@ Abstract
We present a numerical investigation of primordial black hole (PBH) formation from super-horizon curvature perturbations and the subsequent generation and propagation of sound waves, which can serve as a new source of stochastic gravitational wave backgrounds presented in a companion letter. Using the Misner-Sharp formalism with an excision technique, our simulations extend to significantly later times than previous work and indicate that the near-critical perturbations produce a distinct compression wave featuring both overdense and underdense shells, while significantly supercritical perturbations yield only an underdense shell. We also show that a softer equation of state suppresses the formation of compression waves. Furthermore, the comoving thickness of sound shells remains nearly constant during propagation and scales with the Hubble radius at horizon re-entry, thereby serving as a key link between the gravitational-wave peak frequency and PBH mass in the companion letter. These results offer new insights into the dynamics of PBH formation and suggest potential observational signatures of PBHs in the gravitational wave spectrum from associated sound waves.
@@ I Introduction
In the early Universe, matter can be effectively modeled as a fluid medium capable of propagating compression waves, which manifest as sound waves in the small-amplitude regime. During cosmological first-order phase transitions (FOPTs) Mazumdar and White (2019); Hindmarsh et al. (2021); Athron et al. (2024), vacuum bubbles nucleate and expand, perturbing the surrounding fluid. Once they reach terminal velocity Cai and Wang (2021), the bubble walls continuously drive the fluid, generating sound waves with self-similar velocity profiles that depend solely on a dimensionless ratio between radial and temporal coordinates Kurki-Suonio (1985); Kamionkowski et al. (1994); Kurki-Suonio and Laine (1995, 1996). These profiles serve as initial conditions for the sound shell model Hindmarsh (2018); Hindmarsh and Hijazi (2019); Guo et al. (2021), which enables semi-analytic calculations of gravitational wave spectra from the sound waves—one of the dominant sources of gravitational waves during FOPTs Hindmarsh et al. (2014, 2015).
@@
In addition to FOPTs, the formation of primordial black holes (PBHs) Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974) represents another violent process in the early Universe. Gravitational collapse during PBH formation induces intense interactions with ambient fluids, potentially ejecting material outwards as compression waves. In extreme cases, shock waves analogous to supernova ejecta may form, a phenomenon sometimes termed “primordial supernovae” García-Bellido et al. (2021). Collisions between these sound waves are expected to generate gravitational waves as well. Although the sound shell model could be adapted to calculate their spectra, it requires the density and velocity profiles of the sound waves as input. In contrast to the semi-analytically derived self-similar velocity profiles from FOPTs, the configurations of PBH-induced sound waves are governed by nonlinear dynamics of gravitational collapse, necessitating full numerical simulations—the focus of this paper. The gravitational wave spectra derived from these profiles are detailed in a companion letter Zeng et al. (2025).
@@
Beyond gravitational wave generation, PBH-induced compression waves have further cosmological implications. Carr et al. Carr et al. (2021); García-Bellido et al. (2021) noted that shock waves from PBH formation carry substantial energy and exhibit effective temperatures far exceeding that of the ambient plasma. These high-density hotspots could facilitate baryogenesis via electroweak sphaleron transitions Asaka et al. (2004), naturally explaining both the observed baryon-to-photon ratio and the comparable densities of baryonic and dark matter. Although energy conservation arguments suggest that a smaller efficiency factor (thus a lighter PBH mass given the same horizon mass) may produce stronger waves, accurate numerical simulations remain essential for quantifying this energy.
@@
Moreover, the dissipation of sound waves due to photon diffusion Silk (1968) can inject energy into the background, resulting in a spectral distortion in the cosmic microwave background (CMB). This effect was investigated in Refs. Deng et al. (2018); Deng (2020) as a result of the formation of PBHs from primordial bubbles, where the sound waves are generated by the wall-fluid interaction. It was later generalized to any PBH formation mechanisms where perturbations vanish as the spacetime approaches a flat Friedmann-Lemaître-Robertson-Walker (FLRW) Universe asymptotically Deng (2021). In this case, a compensating underdense structure surrounds each PBH and propagates outwards as a sound shell. Therefore, distortions in the CMB spectrum can be used to constrain the abundance of PBHs in general. However, it remains necessary to examine the sound wave generation in other PBH formation mechanisms, such as the collapse of curvature perturbations.
@@
PBH formation from super-horizon curvature perturbations has been extensively studied using numerical simulations, which is the primary focus of this work. In this scenario, the PBH is formed if the amplitude of horizon-scale perturbation exceeds a critical threshold. A key question in this context is whether shock waves and compression waves are generated. In Ref. Hawke and Stewart (2002), the authors used a high-resolution shock-capturing (HRSC) scheme to handle potential shocks. Under sub-horizon perturbations, they detected shock waves in both slightly subcritical and supercritical evolutions, where the energy density varied by several orders of magnitude over a few grid points. Conversely, in Refs. Musco et al. (2005); Polnarev and Musco (2007); Musco et al. (2009), the authors used the Hernandez-Misner formalism Hernandez and Misner (1966) together with an adaptive mesh refinement (AMR) scheme and found no evidence of shock waves under super-horizon perturbations (which is more consistent with inflationary cosmology), even though the perturbation amplitudes were closer to the formation threshold than in the earlier study. They attributed this difference to the different initial conditions used. Although no supersonically propagating shock waves were detected, bounces were observed under subcritical conditions  [^1] In this case, no black hole is formed and only the Misner-Sharp formalism Misner and Sharp (1964) was used. The differences between the two formalisms are discussed in Section II., accompanied by compression waves propagating outwards at the speed of sound. In the supercritical case, however, the formation of compression waves was not reported. We suspect that this omission is due to the insufficient evolution time in their simulations, as compression waves emerge much later than the PBH formation. Another contributing factor may be that the physical quantities were plotted on different null slices, rendering the compression waves less apparent in the figures.
@@
In contrast, Ref. Niemeyer and Jedamzik (1999) observed both bounce and compression waves in slightly supercritical cases using local time-slice plots. They found that as the perturbation amplitude increases, the bounce becomes progressively weaker and eventually disappears because the pressure gradient at the black hole horizon fails to overcome gravitational attraction. However, the above study did not quantify the shape or energy of the compression waves. In contrast to previous work that used the Hernandez-Misner formalism, in this paper, we employ the Misner-Sharp formalism combined with an excision technique to investigate PBH formation and the associated sound waves. We first reproduce the formation of compression waves along with its dynamical profile observed in the previous literature for the subcritical case, and then we demonstrate the generation of sound waves in the supercritical case, analyze their propagation and energy, and discuss their implications for gravitational wave production.
@@
The rest of the paper is organized as follows. In Section II, we describe our numerical framework. Section III presents our simulation results and analyses of the sound waves. Finally, Section IV contains our conclusions and discussion. Throughout this paper, natural units with $c=G=1$ are used.
@@ II Simulation setup
As with most of the other literature on PBH formation, we restrict our attention to the case of spherical symmetry (see Ref. Escrivà (2022) for a review). This simplification is justified by the peak theory Bardeen et al. (1986), which suggests that large peaks tend to be almost spherically symmetric.
The Misner-Sharp Misner and Sharp (1964) and Hernandez-Misner Hernandez and Misner (1966) formalisms both describe the evolution of spherically symmetric relativistic fluids in curved spacetime using a Lagrangian formulation. The Misner-Sharp formalism employs a diagonal metric in which the time coordinate reduces to the familiar FLRW time in the absence of perturbations, commonly referred to as “cosmic time”. This time slice is a space-like hypersurface, which is intuitive. The Misner-Sharp formalism has a well-known drawback for simulating gravitational collapse leading to black hole formation, in that singularities appear shortly after an apparent horizon forms, and the simulation breaks down, precluding the tracking of subsequent evolution (including black hole accretion and sound wave generation). However, this difficulty can be overcome by excising the region where the singularity would develop, as was done in Escrivà (2020). In contrast, the Hernandez-Misner formalism uses an outward null slicing where the time coordinate is called “observer time”. In this coordinate system, it takes infinite coordinate time for an apparent horizon to form, thereby avoiding singularities. Consequently, previous work has mainly used this formalism to follow the PBH evolution and determine their final masses. However, we find that the Hernandez-Misner formalism develops steep gradients in the coordinate space during the later stages of numerical simulations Bloomfield et al. (2015). Resolving these features requires a computationally expensive AMR scheme. Therefore, we adopt the Misner-Sharp formalism with an excision technique for its intuitive cosmic-time framework and capacity to probe post-PBH sound wave generation.
@@ II.1 Misner-Sharp formalism
where $\rho$ denotes the energy density, $p$ the pressure, and $u^{\mu}=A^{-1}\delta^{\mu}_{0}$ the comoving four-velocity of the fluid. Following the notation in Ref. Misner and Sharp (1964), we introduce the following auxiliary variables:
@@
Eq. (5b) is the Hamiltonian constraint, which we use to monitor numerical errors (see Appendix A). The remaining equations govern the temporal evolution of the gravitational system. The boundary conditions are $R(t,r=0)=0$, $U(t,r=0)=0$, $M(t,r=0)=0$, and $D_{r}\rho(t,r=0)=0$ for spherical symmetry, along with an arbitrary nonzero boundary condition for $A$.
@@ II.2 Cosmological setup
Conventional numerical approaches typically evolve the primitive variables $R$, $U$, $M$, and $\rho$ directly. In this work, we need to simulate the system over a much longer time to study the formation and propagation of sound waves. Therefore, to enhance numerical accuracy and stability, we follow Refs. Polnarev et al. (2012); Bloomfield et al. (2015) to factor out the background evolution from the dynamical variables. Specifically, we redefine
@@
To accelerate the numerical evolution, we introduce the logarithmic time variable $\xi\equiv\ln(t/t_{0})$, so that $\partial_{t}=\frac{H}{\alpha}\partial_{\xi}$. Additionally, we define the dimensionless coordinate radius $\bar{r}\equiv r/R_{H}$, where $R_{H}=H^{-1}(t_{0})=t_{0}/\alpha$ is the initial Hubble radius. With these new variables, the equations of motion transform to:
@@ II.3 Initial conditions
For perturbations at super-horizon scales, we use the long-wavelength approximation Salopek and Bond (1990) (which is also called the gradient expansion method) to set up consistent initial conditions. We consider a cosmological perturbation with an initial length scale $r_{m}$ much larger than the Hubble horizon radius $H^{-1}$ and introduce a parameter $\epsilon$ to relate these two scales:
$\epsilon(t)\equiv\frac{H^{-1}(t)}{a(t)r_{m}}=\frac{1}{a(t)H(t)r_{m}}\ll 1.$ (12)
@@
where the curvature perturbation $K(r)$ encodes the deviation from homogeneity. Following Ref. Polnarev and Musco (2007), the Misner-Sharp equations are solved perturbatively:
@@
In the limit $\epsilon\rightarrow 0$, the FLRW solution is recovered. The tilde variables have been computed in Refs. Polnarev and Musco (2007); Musco (2019) and are given by
$\tilde{U}(r)$ $=-\frac{1}{5+3\omega}K(r)r_{m}^{2},$ (15a)
@@
$\tilde{\rho}(r)$ $=\frac{3(1+\omega)}{5+3\omega}\left[K(r)+\frac{r}{3}K^{\prime}(r)\right]r_{m}^{2},$ (15c)
@@
The initial perturbation length scale $r_{m}$ is identified as the location where the compactness function Shibata and Sasaki (1999) (which quantifies the mass excess within a given volume),
@@
with $f(\omega)=3(1+\omega)/(5+3\omega)$. Then $r_{m}$ is related to the curvature perturbation $K(r)$ by $\mathcal{C}^{\prime}(r_{m})=0$, which gives
$K(r_{m})+\frac{r_{m}}{2}K^{\prime}(r_{m})=0.$ (19)
@@
$\delta(t,r)\equiv\frac{1}{V}\int_{0}^{R}4\pi R^{2}\frac{\rho-\rho_{b}}{\rho_{b}}\mathrm{d}R,$ (21)
@@
$\delta(t,r)=\epsilon^{2}(t)\tilde{\delta}(r),$ (22)
with $\tilde{\delta}(r)=f(\omega)K(r)r_{m}^{2}=r_{m}^{2}\mathcal{C}(r)/r^{2}$. The perturbation amplitude is then defined as
$\delta_{m}\equiv\delta(r_{m},t_{m})=\tilde{\delta}(r_{m})=f(\omega)K(r_{m})r_{m}^{2}=\mathcal{C}(r_{m}).$ (23)

### v3

@@ Abstract
We present a numerical investigation of primordial black hole (PBH) formation from super-horizon curvature perturbations and the subsequent generation and propagation of sound waves, which can serve as a new source of stochastic gravitational wave backgrounds (SGWBs) presented in a companion letter. Using the Misner-Sharp formalism with an excision technique, our simulations extend to significantly later times than previous work and indicate that the near-critical perturbations produce a distinct compression wave featuring both overdense and underdense shells, while significantly supercritical perturbations yield only an underdense shell. We also show that a softer equation of state suppresses the formation of compression waves. Furthermore, the comoving thickness of sound shells remains nearly constant during propagation and scales with the Hubble radius at horizon re-entry, thereby serving as a key link between the gravitational-wave peak frequency and PBH mass in the companion letter. These results offer new insights into the dynamics of PBH formation and suggest potential observational signatures of PBHs in the gravitational wave (GW) spectrum from associated sound waves.
@@ I Introduction
In the early Universe, matter can be effectively modeled as a fluid medium capable of propagating compression waves, which manifest as sound waves in the small-amplitude regime. During cosmological first-order phase transitions (FOPTs) Mazumdar:2018dfl; Hindmarsh:2020hop; Athron:2023xlk, vacuum bubbles nucleate and expand, perturbing the surrounding fluid. Once they reach terminal velocity Cai:2020djd, the bubble walls continuously drive the fluid, generating sound waves with self-similar velocity profiles that depend solely on a dimensionless ratio between radial and temporal coordinates Kurki-Suonio:1984zeb; Kamionkowski:1993fg; Kurki-Suonio:1995rrv; Kurki-Suonio:1995yaf. These profiles serve as initial conditions for the sound shell model 1608.04735; 1909.10040; 2007.08537, which enables semi-analytic calculations of gravitational wave (GW) spectra from the sound waves—one of the dominant sources of GWs during FOPTs 1304.2433; 1504.03291.
@@
In addition to FOPTs, the formation of primordial black holes (PBHs) Zeldovich:1967lct; Hawking:1971ei; Carr:1974nx represents another violent process in the early Universe. Gravitational collapse during PBH formation induces intense interactions with ambient fluids, potentially ejecting material outwards as compression waves. In extreme cases, shock waves analogous to supernova ejecta may form, a phenomenon sometimes termed “primordial supernovae” 1904.11482. Collisions between these sound waves are expected to generate GWs as well. Although the sound shell model could be adapted to calculate their spectra, it requires the density and velocity profiles of the sound waves as input. In contrast to the semi-analytically derived self-similar velocity profiles from FOPTs, the configurations of PBH-induced sound waves are governed by nonlinear dynamics of gravitational collapse, necessitating full numerical simulations—the focus of this paper. The GW spectra derived from these profiles are detailed in a companion letter Zeng:2025law.
@@
Beyond GW generation, PBH-induced compression waves have further cosmological implications. Carr et al. 1904.02129; 1904.11482 noted that shock waves from PBH formation carry substantial energy and exhibit effective temperatures far exceeding that of the ambient plasma. These high-density hotspots could facilitate baryogenesis via electroweak sphaleron transitions hep-ph/0310100, naturally explaining both the observed baryon-to-photon ratio and the comparable densities of baryonic and dark matter. Although energy conservation arguments suggest that a smaller efficiency factor (thus a lighter PBH mass given the same horizon mass) may produce stronger waves, accurate numerical simulations remain essential for quantifying this energy.
@@
Moreover, the dissipation of sound waves due to photon diffusion Silk:1967kq can inject energy into the background, resulting in a spectral distortion in the cosmic microwave background (CMB). This effect was investigated in Refs. 1804.10059; 2003.02485 as a result of the formation of PBHs from primordial bubbles, where the sound waves are generated by the wall-fluid interaction. It was later generalized to any PBH formation mechanisms where perturbations vanish as the spacetime approaches a flat Friedmann-Lemaître-Robertson-Walker (FLRW) Universe asymptotically 2106.09817. In this case, a compensating underdense structure surrounds each PBH and propagates outwards as a sound shell. Therefore, distortions in the CMB spectrum can be used to constrain the abundance of PBHs in general. However, it remains necessary to examine the sound wave generation in other PBH formation mechanisms, such as the collapse of curvature perturbations.
@@
PBH formation from super-horizon curvature perturbations has been extensively studied using numerical simulations, which is the primary focus of this work. In this scenario, the PBH is formed if the amplitude of horizon-scale perturbation exceeds a critical threshold. A key question in this context is whether shock waves and compression waves are generated. In Ref. Hawke:2002rf, the authors used a high-resolution shock-capturing (HRSC) scheme to handle potential shocks. Under sub-horizon perturbations, they detected shock waves in both slightly subcritical and supercritical evolutions, where the energy density varied by several orders of magnitude over a few grid points. Conversely, in Refs. gr-qc/0412063; gr-qc/0605122; 0811.1452, the authors used the Hernandez-Misner formalism Hernandez:1966zia together with an adaptive mesh refinement (AMR) scheme and found no evidence of shock waves under super-horizon perturbations (which is more consistent with inflationary cosmology), even though the perturbation amplitudes were closer to the formation threshold than in the earlier study. They attributed this difference to the different initial conditions used. Although no supersonically propagating shock waves were detected, bounces were observed under subcritical conditions  [^1] In this case, no black hole is formed and only the Misner-Sharp formalism Misner:1964je was used. The differences between the two formalisms are discussed in Section II., accompanied by compression waves propagating outwards at the speed of sound. In the supercritical case, however, the formation of compression waves was not reported. We suspect that this omission is due to the insufficient evolution time in their simulations, as compression waves emerge much later than the PBH formation. Another contributing factor may be that the physical quantities were plotted on different null slices, rendering the compression waves less apparent in the figures.
@@
In contrast, Ref. astro-ph/9901292 observed both bounce and compression waves in slightly supercritical cases using local time-slice plots. They found that as the perturbation amplitude increases, the bounce becomes progressively weaker and eventually disappears because the pressure gradient at the black hole horizon fails to overcome gravitational attraction. However, the above study did not quantify the shape or energy of the compression waves. In contrast to previous work that used the Hernandez-Misner formalism, in this paper, we employ the Misner-Sharp formalism combined with an excision technique to investigate PBH formation and the associated sound waves. We first reproduce the formation of compression waves along with its dynamical profile observed in the previous literature for the subcritical case, and then we demonstrate the generation of sound waves in the supercritical case, analyze their propagation and energy, and discuss their implications for GW production.
@@
The rest of the paper is organized as follows. In Section II, we describe our numerical framework. Section III presents our simulation results and analyses of the sound waves. In Section LABEL:sec:GW, we give an estimation of the GW energy spectrum generated by sound shell collisions. Finally, Section LABEL:sec:conclusion contains our conclusions and discussion. Throughout this paper, units with $c=G=1$ are used.
@@ II Simulation setup
As with most of the other literature on PBH formation, we restrict our attention to the case of spherical symmetry (see Ref. 2111.12693 for a review). This simplification is justified by the peak theory Bardeen:1985tr, which suggests that large peaks tend to be almost spherically symmetric.
The Misner-Sharp Misner:1964je and Hernandez-Misner Hernandez:1966zia formalisms both describe the evolution of spherically symmetric relativistic fluids in curved spacetime using a Lagrangian formulation. The Misner-Sharp formalism employs a diagonal metric in which the time coordinate reduces to the familiar FLRW time in the absence of perturbations, commonly referred to as “cosmic time”. This time slice is a space-like hypersurface, which is intuitive. The Misner-Sharp formalism has a well-known drawback for simulating gravitational collapse leading to black hole formation, in that singularities appear shortly after an apparent horizon forms, and the simulation breaks down, precluding the tracking of subsequent evolution (including black hole accretion and sound wave generation). However, this difficulty can be overcome by excising the region where the singularity would develop, as was done in Escriva:2019nsa. In contrast, the Hernandez-Misner formalism uses an outward null slicing where the time coordinate is called “observer time”. In this coordinate system, it takes infinite coordinate time for an apparent horizon to form, thereby avoiding singularities. Consequently, previous work has mainly used this formalism to follow the PBH evolution and determine their final masses. However, we find that the Hernandez-Misner formalism develops steep gradients in the coordinate space during the later stages of numerical simulations Bloomfield:2015ila. Resolving these features requires a computationally expensive AMR scheme. Therefore, we adopt the Misner-Sharp formalism with an excision technique for its intuitive cosmic-time framework and capacity to probe post-PBH sound wave generation.
@@ II.1 Misner-Sharp formalism
where $\rho$ denotes the energy density, $p$ the pressure, and $u^{\mu}=A^{-1}\delta^{\mu}_{0}$ the comoving four-velocity of the fluid. Following the notation in Ref. Misner:1964je, we introduce the following auxiliary variables:
@@
Eq. (5b) is the Hamiltonian constraint, which we use to monitor numerical errors (see Appendix LABEL:app:convergence). The remaining equations govern the temporal evolution of the gravitational system. The boundary conditions are $R(t,r=0)=0$, $U(t,r=0)=0$, $M(t,r=0)=0$, and $D_{r}\rho(t,r=0)=0$ for spherical symmetry, along with an arbitrary nonzero boundary condition for $A$.
@@ II.2 Cosmological setup
Conventional numerical approaches typically evolve the primitive variables $R$, $U$, $M$, and $\rho$ directly. In this work, we need to simulate the system over a much longer time to study the formation and propagation of sound waves. Therefore, to enhance numerical accuracy and stability, we follow Refs. Polnarev:2012bi; Bloomfield:2015ila to factor out the background evolution from the dynamical variables. Specifically, we redefine
@@
To accelerate the numerical evolution, we introduce the logarithmic time variable $\xi\equiv\ln(t/t_{0})$, so that $\partial_{t}=\frac{H}{\alpha}\partial_{\xi}$. Additionally, we define the dimensionless coordinate radius $\bar{r}\equiv r/R_{H}$, where $R_{H}=H^{-1}(t_{0})=t_{0}/\alpha$ is the initial Hubble radius. In this case, all quantities are dimensionless and our plots are independent of the choice of $R_{H}$. With these new variables, the equations of motion transform to:
@@ II.3 Initial conditions
For perturbations at super-horizon scales, we use the long-wavelength approximation Salopek:1990jq (which is also called the gradient expansion method) to set up consistent initial conditions. We consider a cosmological perturbation with an initial length scale $r_{i}$ much larger than the Hubble horizon radius $H^{-1}$ and introduce a parameter $\epsilon$ to relate these two scales:
$\epsilon(t)\equiv\frac{H^{-1}(t)}{a(t)r_{i}}=\frac{1}{a(t)H(t)r_{i}}\ll 1.$ (12)
@@
where the curvature perturbation $K(r)$ encodes the deviation from homogeneity. Following Ref. gr-qc/0605122, the Misner-Sharp equations are solved perturbatively:
@@
In the limit $\epsilon\rightarrow 0$, the FLRW solution is recovered. The tilde variables have been computed in Refs. gr-qc/0605122; 1809.02127 and are given by
$\tilde{U}(r)$ $=-\frac{1}{5+3\omega}K(r)r_{i}^{2},$ (15a)
@@
$\tilde{\rho}(r)$ $=\frac{3(1+\omega)}{5+3\omega}\left[K(r)+\frac{r}{3}K^{\prime}(r)\right]r_{i}^{2},$ (15c)
@@
The initial perturbation length scale $r_{i}$ is identified as the location where the compactness function gr-qc/9905064 (which quantifies the mass excess within a given volume),
@@
with $f(\omega)=3(1+\omega)/(5+3\omega)$. Then $r_{i}$ is related to the curvature perturbation $K(r)$ by $\mathcal{C}^{\prime}(r_{i})=0$, which gives
$K(r_{i})+\frac{r_{i}}{2}K^{\prime}(r_{i})=0.$ (19)
@@
$\bar{\delta}(t,r)\equiv\frac{1}{V}\int_{0}^{r}4\pi R^{2}\frac{\rho-\rho_{b}}{\rho_{b}}\mathrm{d}R,$ (21)
@@
$\bar{\delta}(t,r)=\epsilon^{2}(t)\tilde{\delta}(r),$ (22)
with $\tilde{\delta}(r)=f(\omega)K(r)r_{i}^{2}=r_{i}^{2}\mathcal{C}(r)/r^{2}$. The perturbation amplitude is then defined as
$\delta_{m}\equiv\bar{\delta}(r_{i},t_{m})=\tilde{\delta}(r_{i})=f(\omega)K(r_{i})r_{i}^{2}=\mathcal{C}(r_{i}).$ (23)